[figure: 9.7 cm]

ELEMENTARY EXCITATIONS IN HEISENBERG ULTRA-THIN FILMS WITH INHOMOGENEUS SPATIAL DISTRIBUTION OF

MAGNETIZATION

Z. Onyszkiewicz and R. Łosiński

Physics Department, Adam Mickiewicz University, Umultowska 85, 61-614 Poznań, Poland

Taking advantage of the generalised equilibrium reduced density operator^{1} as well as the normalised high-density expansion method for the temperature matrix casual Green function^{2}, the elementary excitations fot the Heisenberg (S = \frac{1}{2}) ultra-thin films with homogeneus and inhomogeneus spatial distribution of magnetization were determined. Moreover, the temperature dependence of elementary excitations in such ultra-thin films of cubic structure was obtained. The effect of non-magnetic substrate on the spectrum of elementary excitations was analysed, by assuming a linear dependence of the exchange parameters in the monoatomic film adjacent to the substrate on the appropriate component of the film deformation tensor.

1. Z. Onyszkiewicz, Physics Lett. A 163, 450 (1992).

2. Z. Onyszkiewicz, Physica B 151, 462 (1988).

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : Z. Onyszkiewicz

Address for correspondence : Physics Department, Adam Mickiewicz University, Umultowska 85, 61-614 Poznań, Poland

Email address : zbigonys@main.amu.edu.pl